OpenStudy (anonymous):

what are the factors of x^2-5x+6

OpenStudy (anonymous):

factored that becomes (x-6)(x+1)

OpenStudy (anonymous):

How do you solve it it step ?

OpenStudy (anonymous):

oh wait I was actually wrong its (x+3)(x+2)

OpenStudy (anonymous):

which factors of 6 will add up to 5 and the product is 6? 3 and 2. 3+2 = 5 and 3x2=6 so actually i was wrong again lol (sorry) its actually (x-3)(x-2)

OpenStudy (anonymous):

-3x-2=6 also -3-2=-5

OpenStudy (anonymous):

would you please solve step by step

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Factoring x2-5x+6 The first term is, x2 its coefficient is 1 . The middle term is, -5x its coefficient is -5 . The last term, "the constant", is +6 Step-1 : Multiply the coefficient of the first term by the constant 1 • 6 = 6 Step-2 : Find two factors of 6 whose sum equals the coefficient of the middle term, which is -5 . -6 + -1 = -7 -3 + -2 = -5 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -3 and -2 x2 - 3x - 2x - 6 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-3) Add up the last 2 terms, pulling out common factors : 2 • (x-3) Now add up the four terms of step 3 : (x-2) • (x-3) Which is the desired factorization Final result : (x - 2) • (x - 3)

OpenStudy (anonymous):

please solve in formation